stanley has $1.45 in dimes and quarters. If he had 10 coins total, how many of each coin does he have.
Answer by Alan3354(69443)

stanley has $1.45 in dimes and quarters. If he had 10 coins total, how many of each coin does he have.
------------------
If all 10 coins were dimes, he'd have $1.00
Each dime replaced by a quarter adds 15 cents.
---
145 - 100 = 45
45/15 = 3
--> 3 quarters, 7 dimes
=============================
The hard way:
D + Q = 10
10D + 25Q = 145
------------
etc.